OpenStudy (anonymous):

A. write the expression 81^-3/4 in radical notation. B. evaluate the radical expression

OpenStudy (anonymous):

im lost pl help me

OpenStudy (lukecrayonz):

1/27

OpenStudy (lukecrayonz):

and decimal form is 0.037...

OpenStudy (anonymous):

ok 0.037 is

OpenStudy (lukecrayonz):

@Directrix he started messaging me and I explained:)

Directrix (directrix):

I would like to know the steps for my own learning.

OpenStudy (dan815):

lol

OpenStudy (lukecrayonz):

Sure! :P 81^-3/4 you can use the laws of exponents to rewrite it A positive rational exponent is b^m/n= nth root (b^m) A negative exponent would be b^-m/n = 1/b^m/n (when you have a negative exponent, you can rewrite it over one so it's positive)

OpenStudy (anonymous):

84^3/4 is 27, so 1/27 got my anwser

OpenStudy (dan815):

beautiful

OpenStudy (anonymous):

a is 27 and b 1/27

OpenStudy (dan815):

noo

OpenStudy (anonymous):

ok im dum

Directrix (directrix):

How did you get 1/27, the explicit steps for that is what I am asking. @Lukecrayonz And, what about this part, " write the expression 81^-3/4 in radical notation." What is the answer for that?

OpenStudy (dan815):

oh i just noticed u changed ur 81 to 84!! how did u get this

OpenStudy (anonymous):

no it 81

OpenStudy (anonymous):

im sorry

OpenStudy (skullpatrol):

Did you follow my last explanation?

Directrix (directrix):

I think this is the same question: http://openstudy.com/study#/updates/527226d5e4b077aaee84128b